# Contextual Document Embeddings

### v3

@@ 4.1 Contextual Training with Adversarial Contrastive Learning
Computationally, the inner term can be implemented as a single batch and computed efficiently without the need for separate hard negatives ($\mathcal{H}$). Ideally we want groups that are as challenging as possible. Zhang & Stratos (2021) show that increasing the partition term improves the contrastive approximation to the maximum likelihood the gradient. We can formalize the search for the most difficult configuration of batches as an optimization problem:
$\max_{({\cal B}^{1},\ldots{\cal B}^{B})}\sum_{b}\sum_{\begin{subarray}{c}(d,q)\in{\cal B}^{b}\\ (d^{\prime},q^{\prime})\in{\cal B}^{b}\end{subarray}}f(d,q^{\prime})+f(d^{\prime},q)=\max_{({\cal B}^{1},\ldots{\cal B}^{B})}\sum_{b}\sum_{\begin{subarray}{c}(d,q)\in{\cal B}^{b}\\ (d^{\prime},q^{\prime})\in{\cal B}^{b}\end{subarray}}\phi(d)\cdot\psi(q^{\prime})+\phi(d^{\prime})\cdot\psi(q)$ (2)
Solving this combinatorial objective exactly is intractable, but we can treat approximate a solution using clustering. We first move from a maximization to a minimization by replacing the two dot products with L2, i.e. $m((d,q),(d^{\prime},q^{\prime}))=||\phi(d)-\psi(q^{\prime})||+||\phi(^{\prime}d)-\psi(q)||$ which is equivalent for normalized embeddings. We then note that treated as symmetric pairs, this term obeys the triangle inequality for any other pair $m$ i.e:
@@
For a known size $B$, this defines an asymmetric K-Means clustering problem. A solution can be efficiently computed using extremely fast Euclidean K-Means packages be treating each data point as two separate vectors $\phi(d)\oplus\psi(q)$ and $\psi(q)\oplus\phi(d)$ where $\oplus$ is concatenation.
@@
Since clustering is performed before training, we do not have dense encoders for $\phi$ and $\psi$ when constructing the groups. Borrowing methods from hard-negative mining (Robinson et al., 2021) we can replace the $\phi$ and $\psi$ with a simpler embedding model when constructing groups. We experiment with a sparse vector representation and with pretrained dense representations, settling on GTR (Ni et al., 2021), a popular and generic text embedding model.
@@ 4.2 Contextual Document Embedding (CDE)
Contextualization can also be added directly to the architnhlltivvffrrhgcfvbjgnurgkhbclnlvecture. Taking inspiration from sparse vector retrieval which uses corpus statistics to determine the form of the embedding, we modify the encoders to have access to the corpus itself, i.e. $\phi(d;{\cal D})$ and $\psi(d;{\cal D})$. This effectively augments the biencoder model to give it the ability to contextualize documents directly.

### v4

@@ 4.1 Contextual Training with Adversarial Contrastive Learning
Computationally, the inner term can be implemented as a single batch and computed efficiently without the need for separate hard negatives ($\mathcal{H}$). Ideally we want groups that are as challenging as possible. Zhang & Stratos (2021) show that increasing the partition term improves the contrastive approximation to the maximum likelihood of the gradient. We can formalize this search for the most difficult configuration of batches as an optimization problem:
$\max_{({\cal B}^{1},\ldots{\cal B}^{B})}\sum_{b}\sum_{\begin{subarray}{c}(d,q)\in{\cal B}^{b}\\ (d^{\prime},q^{\prime})\in{\cal B}^{b}\end{subarray}}\left(f(d,q^{\prime})+f(d^{\prime},q)\right)=\max_{({\cal B}^{1},\ldots{\cal B}^{B})}\sum_{b}\sum_{\begin{subarray}{c}(d,q)\in{\cal B}^{b}\\ (d^{\prime},q^{\prime})\in{\cal B}^{b}\end{subarray}}\left(\phi(d)\cdot\psi(q^{\prime})+\phi(d^{\prime})\cdot\psi(q)\right)$ (2)
Solving this combinatorial objective exactly is intractable, but we can approximate a solution using clustering. We first move from a maximization to a minimization by replacing the two dot products with L2 distance $m((d,q),(d^{\prime},q^{\prime}))=||\phi(d)-\psi(q^{\prime})||+||\phi(^{\prime}d)-\psi(q)||$ (which is equivalent for normalized embeddings). We then note when that treated as symmetric pairs, this term obeys the triangle inequality for any other pair $m$:
@@
For known $B$, this search defines an asymmetric K-Means clustering problem. A solution can be efficiently computed using extremely fast Euclidean K-Means packages be treating each data point as two separate vectors $\phi(d)\oplus\psi(q)$ and $\psi(q)\oplus\phi(d)$, where $\oplus$ is concatenation.
@@
Since clustering is performed before training, we do not have dense encoders $\phi$ and $\psi$ when constructing the groups. Borrowing methods from hard-negative mining (Robinson et al., 2021) we can replace the $\phi$ and $\psi$ with a simpler embedding model when constructing groups. We experiment with a sparse vector representation and with pretrained dense representations, settling on GTR (Ni et al., 2021), a popular and generic text embedding model.
@@ 4.2 Contextual Document Embedding (CDE)
Contextualization can also be added directly to the architecture. Taking inspiration from sparse vector retrieval which uses corpus statistics to determine the form of the embedding, we modify the encoders to have access to the corpus itself, i.e. $\phi(d;{\cal D})$ and $\psi(d;{\cal D})$. This effectively augments the biencoder model to give it the ability to contextualize documents directly.